OpenStudy (anonymous):

The probability that rose will complete a free through basketball shot is 1/3. If she makes 5 shots, what is the probability of her completing at least one shot?

OpenStudy (perl):

this is a binomial problem. we want completing at least one shot. Let X = # of shots in 5 chances . Then X can equal to 0,1,2,3,4,5 We want P( X >= 1 ) or P (X = 1,2,3,4,5)

OpenStudy (perl):

each shot has probability of success 1/3 and failure 2/3 . it is easier to do the complement P ( X > = 1 ) = 1 - P(X <1 ) = 1 - P( X = 0 ) = 1 - ( 2/3)^5

OpenStudy (perl):

.8683127572